A new car is purchased for 21100 dollars. The value of the car depreciates at 10.5% per year. What will the value of the car be, to the nearest cent, after 6 years?


1
Expert's answer
2021-03-29T17:00:19-0400

Answer

Value of car after 6 year

"V(t) =21100(1-\frac{10.5}{100}) ^6\\V(t) =10844.78"

So value of the car after 6 years will be 10844.78$.